OpenStudy (anonymous):

use the properties of limits to help decide whether the limit exists. if limit exist find it's value... lim -4x^6+ 3x^3 + 6x / x> infinity 3x^2 - 7x +2

sam (.sam.):

I assumed \[\huge \lim_{x \rightarrow \infty} \frac{-4x^6+ 3x^3 + 6x}{ 3x^2 - 7x +2}\]

OpenStudy (anonymous):

Yes

sam (.sam.):

ok, we have the highest degree at the numerator, so when x approaches infiinity (means x is getting larger values), the whole fraction gets very large because the power is higher at numerator. and the highest power is negative, so you will get \[- \infty\]

sam (.sam.):

That limit is not exist

OpenStudy (anonymous):

Awesome! Thank you so much!